# Asymptotic computations of tropical refined invariants in genus 0 and 1

### v1

@@ 4.1. The case of Hirzebruch surfaces
We consider words over the alphabet $\{\mathsf{f},\mathsf{e},\mathsf{b}_{j},\mathsf{t}_{j}\}_{j\in\mathbb{N}}$. The letters used stand for “floor”, “edge/elevator”, “bottom end” and “top end”. The indices of the letters refer to the number of floors they skip. We first explain how to get a word $W(\mathcal{D})$ from a genus 00 marked diagram $\mathcal{D}$ whose floors are totally ordered. Let $aE+bF\in H_{2}(\mathbb{F}_{\delta},\mathbb{Z})$ be the class of the diagram. The floors of $\mathcal{D}$ are labelled from $1$ to $a$. The letters of the word $W(\mathcal{D})$ are in ordered correspondence with the marked points of $\mathcal{D}$ with the following rule:
@@
that associate to a sentence in $\mathcal{S}$ the sum of the codegrees of its words, the length of the words $\mathsf{S}_{0}$ and $\mathsf{S}_{j}^{(k)}$ (maybe 00), and the sum of their lengths. For $n\geqslant 0$ we denote by $\mathcal{S}(n)$ the set of sentences with total length $n$. Lemma 4.4 asserts that choosing a word in $\mathscr{W}_{aE+bF}$ having codegree at most $i$ and with $a,b$ large enough amounts to choose :
@@ 4.2. The case of $h$-transverse toric surfaces
Compared with the Hirzebruch case, the marked floor diagrams are modified by incorporating the data $(L,R)$, i.e. assigning a pair of integers called sloping pair to each floor. According to [BJP22, Section 3], the codegree coming from the sloping pairs is
@@
We denote by $(L,R)$ the tuple of sloping pairs of $\mathsf{W}$. We first notice that the tuple $L$ (and similarly for $R$) differs from the unique tuple of increasing slopes by a finite number of transpositions that switches two consecutive elements. Indeed, this is true for the tuple of codegree 0 since in that case this tuple is increasing. If we consider a tuple of positive codegree then there is a consecutive pair that forms an inversion; if not, the tuple would be increasing. Then, switching both members of the inversion decreases the codegree, and we conclude by induction. Each transposition switching consecutive elements increases the codegree by at least $1$, so that if $\mathrm{codeg}(\mathsf{W})\leqslant i$ then $L$ differs from the increasing tuple by at most $i$ transpositions.
@@ 5.1. The case of Hirzebruch surfaces
Assume $b>i+2M$ and let $\mathcal{D}$ be a diagram in the class $aE+bF$ with $\mathrm{codeg}(\mathcal{D})\leqslant i$.
@@
$\int_{\mathcal{L}}\ell_{m}^{(r)}\mathrm{d}\nu_{\mathcal{L}}=\frac{1}{p(x)^{2}}\sum_{\mathbf{l}\in\mathcal{L}}l_{m}^{(r)}\prod_{j,k}x^{jl_{j}^{(k)}}=\frac{1}{p(x)^{2}}\left(\sum_{l_{m}^{(r)}=0}^{\infty}l_{m}^{(r)}x^{ml_{m}^{(r)}}\right)\prod_{(j,k)\neq(m,r)}\left(\sum_{l_{j}^{(k)}=0}^{\infty}x^{jl_{j}^{(k)}}\right).$
@@
In fact, this method, which is an analog of Fubini’s theorem, works for computing the integral of any monomial in the $\ell_{j}^{(k)}$: the integral of a monomial is equal to the product of integrals over each of the variables appering in the monomial. Hence, it reduces down to the computation of the sums $\sum_{\alpha=0}^{\infty}\alpha^{r}y^{\alpha}$.
@@ 5.1.5. Computation of the asymptotic refined invariant
$\circ$ First step: expression over $\mathcal{S}(b)\times\mathcal{S}(b+\delta a)$. By Lemma 5.9, the function giving the sum of multiplicities for insertion of a marking between the floors $m$ and $m+1$ is
$\mathrm{pos}_{m}\left(\frac{\widetilde{\omega}_{m}-1}{2}-d_{m}\right)=\left\{\begin{array}[]{ll}(\rho_{1}^{*}\ell_{m}+2)\left(\dfrac{\omega_{m}-\rho_{1}^{*}\phi_{m}[b]-1}{2}-\langle m\rangle\right)&\text{ if }m\leqslant i,\\ \omega_{m}-1&\text{ if }i<m<a-i\\ (\rho_{2}^{*}\ell_{a-m}+2)\left(\dfrac{\omega_{m}-\rho_{2}^{*}\phi_{a-m}[b+\delta a]-1}{2}-\langle a-m\rangle\right)&\text{ if }m\geqslant a-i,\end{array}\right.$
where in the first (resp. last) row, functions are pull-back from $\mathcal{S}(b)$ (resp. $\mathcal{S}(b+\delta a)$). For each value of $m$, we now need to integrate the above function, and then sum over $1\leqslant m\leqslant a$.
@@
$\int_{\mathcal{S}(b)\times\mathcal{S}(b+\delta a)}(\rho_{1}^{*}\ell_{m}+2)\left(\dfrac{\omega_{m}-\rho_{1}^{*}\phi_{m}[b]-1}{2}-\langle m\rangle\right)=\int_{\mathcal{S}(b)}(\ell_{m}+2)\left(\dfrac{\omega_{m}-\phi_{m}[b]-1}{2}-\langle m\rangle\right).$
@@
$\int_{\mathcal{S}(b)\times\mathcal{S}(b+\delta a)}\mathrm{pos}_{m}\left(\frac{\widetilde{\omega}_{m}-1}{2}-d_{m}\right)=\omega_{m}-1+(\delta-2)m\frac{x^{m}}{1-x^{m}}-2\sum_{j=m}^{\infty}\langle j\rangle-2\frac{x^{m}}{(1-x^{m})^{2}}.$
@@
$\int_{\mathcal{S}(b)\times\mathcal{S}(b+\delta a)}\mathrm{pos}_{m}\left(\frac{\widetilde{\omega}_{m}-1}{2}-d_{m}\right)=\omega_{m}-1-(\delta+2)m^{\prime}\frac{x^{m^{\prime}}}{1-x^{m^{\prime}}}-2\sum_{j=m^{\prime}}^{\infty}\langle j\rangle-2\frac{x^{m^{\prime}}}{(1-x^{m^{\prime}})^{2}}.$
@@
The computations made in the Hirzebruch case remain valid with two differences. First, we now need to take into account the sloping pairs to the floors. The marked diagrams of genus 0 and codegree at most $i$ are in bijection with a subset of $\mathcal{S}(b^{\mathrm{bot}})\times\mathcal{S}(b^{\mathrm{top}})\times\mathcal{P}^{\chi-4}$, where $\chi$ is the number of corners of the polygon and $\mathcal{P}$, defined in Section 4.2.2, encodes the default of increasingness of the slopes. Second, the self-intersection of the divisors corresponding to the top and bottom horizontal sides, equal to $\delta$ and $-\delta$ in the Hirzebruch case, are not opposite anymore, see Lemma 2.5.
@@
$\sum_{\begin{subarray}{c}\mathfrak{P}=(\mathfrak{p}_{c})_{c}\\ \mathrm{codeg}(\mathfrak{p}_{c})\leqslant i\end{subarray}}x^{\sum\mathrm{codeg}(\mathfrak{p}_{c})}p(x)^{4}\left[\left(g_{\max}-\sum_{c}\mathrm{codeg}(\mathfrak{p}_{c})\right)-(12-\delta_{\mathrm{top}}-\delta_{\mathrm{bot}})E_{2}(x)\right].$

### v2

@@ 4.1. The case of Hirzebruch surfaces
We consider words over the following alphabet : $\{\mathsf{f},\mathsf{e},\mathsf{b}_{j},\mathsf{t}_{j}\}_{j\in\mathbb{N}}$. The letters used stand for “floor”, “edge/elevator”, “bottom end” and “top end”. The indices of the letters refer to the number of floors they skip. We first explain how to get a word $W(\mathcal{D})$ from a genus 00 marked diagram $\mathcal{D}$ whose floors are totally ordered. Let $aE+bF\in H_{2}(\mathbb{F}_{\delta},\mathbb{Z})$ be the class of the diagram. The floors of $\mathcal{D}$ are labelled from $1$ to $a$. The letters of the word $W(\mathcal{D})$ are in ordered correspondence with the marked points of $\mathcal{D}$ with the following rule:
@@
that associate to a sentence in $\mathcal{S}$ the sum of the codegrees of its words, the length of the words $\mathsf{S}_{0}$ and $\mathsf{S}_{j}^{(k)}$ (maybe 00), and the sum of their lengths. For $n\geqslant 0$ we denote by $\mathcal{S}(n)$ the set of sentences with total length $n$.
Lemma 4.4 asserts that choosing a word in $\mathscr{W}_{aE+bF}$ having codegree at most $i$ and with $a,b$ large enough amounts to choose :
@@ 4.2. The case of $h$-transverse toric surfaces
The main difference with the Hirzebruch case, is that marked floor diagrams are modified by incorporating the data $(L,R)$, i.e. assigning a pair of integers called sloping pair to each floor. According to [BJP22, Section 3], the codegree coming from the sloping pairs is
@@
We denote by $(L,R)$ the tuple of sloping pairs of $\mathsf{W}$. We first notice that the tuple $L$ (and similarly for $R$) differs from the unique tuple of increasing slopes by a finite number of transpositions that switches two consecutive elements. Indeed, this is true for the tuple of codegree 0 since in that case this tuple is increasing.
If we consider a tuple of positive codegree then there is a consecutive pair that forms an inversion; if not, the tuple would be increasing. Then, switching both members of the inversion decreases the codegree, and we conclude by induction.
Each transposition switching consecutive elements increases the codegree by at least $1$, so that if $\mathrm{codeg}(\mathsf{W})\leqslant i$ then $L$ differs from the increasing tuple by at most $i$ transpositions.
@@ 5.1. The case of Hirzebruch surfaces
Assume $b>i+2M$ and let $\mathcal{D}$ be a floor diagram in the class $aE+bF$ with $\mathrm{codeg}(\mathcal{D})\leqslant i$.
@@ 3. $(iii)$
$\int_{\mathcal{L}}\ell_{m}^{(r)}\mathrm{d}\nu_{\mathcal{L}}$ $=\frac{1}{p(x)^{2}}\sum_{\mathbf{l}\in\mathcal{L}}l_{m}^{(r)}\prod_{j,k}x^{jl_{j}^{(k)}}$
$=\frac{1}{p(x)^{2}}\left(\sum_{l_{m}^{(r)}=0}^{\infty}l_{m}^{(r)}x^{ml_{m}^{(r)}}\right)\prod_{(j,k)\neq(m,r)}\left(\sum_{l_{j}^{(k)}=0}^{\infty}x^{jl_{j}^{(k)}}\right).$
@@
In fact, this method, which is an analog of Fubini’s theorem, works for computing the integral of any monomial in the $\ell_{j}^{(k)}$: the integral of a monomial is equal to the product of integrals over each of the variables appearing in the monomial. Hence, it reduces down to the computation of the sums $\sum_{\alpha=0}^{\infty}\alpha^{r}y^{\alpha}$.
@@ 5.1.5. Computation of the asymptotic refined invariant
$\circ$ First step: expression over $\mathcal{S}(b)\times\mathcal{S}(b+\delta a)$. By Lemma 5.9, the function giving the sum of multiplicities for insertion of a marking between the floors $m$ and $m+1$ is given by $\mathrm{pos}_{m}\left(\frac{\widetilde{\omega}_{m}-1}{2}-d_{m}\right)$ and has the following values:
$\left\{\begin{array}[]{ll}(\rho_{1}^{*}\ell_{m}+2)\left(\dfrac{\omega_{m}-\rho_{1}^{*}\phi_{m}[b]-1}{2}-\langle m\rangle\right)&\text{ if }m\leqslant i,\\ \omega_{m}-1&\text{ if }i<m<a-i\\ (\rho_{2}^{*}\ell_{a-m}+2)\left(\dfrac{\omega_{m}-\rho_{2}^{*}\phi_{a-m}[b+\delta a]-1}{2}-\langle a-m\rangle\right)&\text{ if }m\geqslant a-i,\end{array}\right.$
where in the first (resp. last) row, functions are pull-back from $\mathcal{S}(b)$ (resp. $\mathcal{S}(b+\delta a)$). For each value of $m$, we now need to integrate the above function, and then sum over $11\leqslant m\leqslant a-$ 1.
@@
$\int_{\mathcal{S}(b)\times\mathcal{S}(b+\delta a)}(\rho_{1}^{*}\ell_{m}+2)\left(\dfrac{\omega_{m}-\rho_{1}^{*}\phi_{m}[b]-1}{2}-\langle m\rangle\right)$
$=$ $\int_{\mathcal{S}(b)}(\ell_{m}+2)\left(\dfrac{\omega_{m}-\phi_{m}[b]-1}{2}-\langle m\rangle\right).$
@@
$\ \int_{\mathcal{S}(b)\times\mathcal{S}(b+\delta a)}\mathrm{pos}_{m}\left(\frac{\widetilde{\omega}_{m}-1}{2}-d_{m}\right)$
$=$ $\ \omega_{m}-1+(\delta-2)m\frac{x^{m}}{1-x^{m}}-2\sum_{j=m}^{\infty}\langle j\rangle-2\frac{x^{m}}{(1-x^{m})^{2}}.$
@@
$\ \int_{\mathcal{S}(b)\times\mathcal{S}(b+\delta a)}\mathrm{pos}_{m}\left(\frac{\widetilde{\omega}_{m}-1}{2}-d_{m}\right)$
$=$ $\ \omega_{m}-1-(\delta+2)m^{\prime}\frac{x^{m^{\prime}}}{1-x^{m^{\prime}}}-2\sum_{j=m^{\prime}}^{\infty}\langle j\rangle-2\frac{x^{m^{\prime}}}{(1-x^{m^{\prime}})^{2}}.$
@@
The computations made in the Hirzebruch case remain valid with two differences. First, we now need to take into account the sloping pairs of the floors. The marked diagrams of genus 0 and codegree at most $i$ are in bijection with a subset of $\mathcal{S}(b^{\mathrm{bot}})\times\mathcal{S}(b^{\mathrm{top}})\times\mathcal{P}^{\chi-4}$, where $\chi$ is the number of corners of the polygon and $\mathcal{P}$, defined in Section 4.2.2, encodes the default of growth of the slopes. Second, the self-intersection of the divisors corresponding to the top and bottom horizontal sides, equal to $\delta$ and $-\delta$ in the Hirzebruch case, are not opposite anymore, see Lemma 2.5.
@@
$\sum_{\begin{subarray}{c}\mathfrak{P}=(\mathfrak{p}_{c})_{c}\\ \mathrm{codeg}(\mathfrak{p}_{c})\leqslant i\end{subarray}}x^{\sum\mathrm{codeg}(\mathfrak{p}_{c})}p(x)^{4}\left[\left(g_{\max}-\sum_{c}\mathrm{codeg}(\mathfrak{p}_{c})\right)-(12-\delta_{\mathrm{top}}-\delta_{\mathrm{bot}})E_{2}(x)\right]$